The mean grade in a Physics examination was 75 and the standard deviation was 5. Determine the standard scores (z-value) corresponding to a grade of 85.

1
Expert's answer
2021-11-10T18:27:35-0500
"Z=\dfrac{X-\mu}{\sigma}=\dfrac{85-75}{5}=2"

The standard scores (z-value) corresponding to a grade of 85 is "Z=2."